# Approximation in Hölder spaces

### v3

@@ 2. Bounded support approximation
Let $f\in\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ and by Theorem 1.2 find $g\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(\mathbb{R}^{n},Y)$ with bounded support such that $\|f-g\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}<\varepsilon.$ Let $r>0$ and $0\leq\eta_{r}\in C^{\infty}_{c}(B(0,r))$ be a standard real-valued smooth bump in $\mathbb{R}^{n}$ with $\int_{\mathbb{R}^{n}}\eta_{r}=1,$ and we define $h_{r}:=g*\eta_{r},$ where the integral is understood in the Bochner sense. Notice that $h_{r}$ is compactly supported, as $g$ is, and smooth $h_{r}\in C^{\infty}(\mathbb{R}^{n},Y)$ as a smooth mollification. Provided that $|u|\leq\delta,$ there holds uniformly in $r$ that
$|h_{r}(x+u)-h_{r}(x)|=\Big{|}\int_{\mathbb{R}^{n}}\eta_{r}(y)(g(x+u+y)+g(x+y))\,\mathrm{d}y\Big{|}\leq\|g\|_{\dot{C}^{0,\omega}_{\delta}}\omega(|u|),$
and hence we find $\delta>0$ so that, uniformly in $r,$ we have $\|g-h_{r}\|_{\dot{C}^{0,\omega}_{\delta}}\leq\|g\|_{\dot{C}^{0,\omega}_{\delta}}+\|h_{r}\|_{\dot{C}^{0,\omega}_{\delta}}\leq\varepsilon.$ To deal with the scales $\geq\delta$ we argue as follows. Observe that $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ is uniformly continuous, by $\lim_{t\to 0}\omega(t)=0,$ and thus the approximation $h_{r}:=g*\eta_{r}$ converges uniformly to $g,$ as $r\to 0.$ We let $r=r(\delta)$ be so small that $\|g-h_{r}\|_{\infty}\leq\frac{1}{2}\varepsilon\omega(\delta).$ Then, for $|x-y|>\delta,$ there holds that
$\|(g-h_{r})(x)-(g-h_{r})(y)\|\leq 2\|g-h_{r}\|_{\infty}\leq\varepsilon\omega(\delta)\leq\varepsilon\omega(|x-y|),$
@@
For the reverse inclusion ``","```$\supset$",` ` ⊃ " , simply notice that by the Mean Value Inequality, $C^{\infty}_{c}(\mathbb{R}^{n},Y)\subset\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$. We remind that this inequality holds for smooth $Y$-valued functions, as a consequence of the Hahn-Banach theorem on $Y.$ Also, by (3) there holds that $\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)\subset\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y).$ As $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ is closed with respect to $\dot{C}^{0,\omega}$ limits, by Remark 2.1, we are done. ∎
@@
For a subset $G$ of $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y),$ we write $G\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ (with ``$u"$ for uniformly) provided that $G$ is equibounded in the $\dot{C}^{0,\omega}$ seminorm, and that for all $\varepsilon>0,$ there exists $t>0$ such that if $|x-y|<t$ or $|x|>t^{-1}$ or $|y|>t^{-1},$ then
@@
We first prove that the norms are equivalent. So fix an arbitrary cube $Q\in\mathcal{Q}$ and estimate
@@
For the other direction, let $x,z\in\mathbb{R}^{n}$ be arbitrary and let $Q_{0}\subset\mathbb{R}^{n}$ be a cube containing both $x$ and $y$ such that $\ell(Q_{0})=|x-y|.$ Consider the dyadic descendants of $Q_{0}$ (achieved by iteratively halving the sides) and for every $k\in\mathbb{N}\cup\{0\}$ let $Q_{k}(x)$ be the descendant of $Q_{0}$ of sidelength $2^{-k}\ell(Q_{0})$ that contains the point $x.$ Similarly, we define $Q_{k}(y)$ for each $k.$ By the continuity of $f:\mathbb{R}^{n}\to Y,$ we write
@@
Both sums are estimated identically, so consider the first:
@@
and we continue the bound, using that $\omega$ is non-decreasing in the first bound, with
$\sum_{k=0}^{\infty}\omega(\ell(Q_{k}(x)))=\sum_{k=0}^{\infty}\ell(Q_{k}(x))\frac{\omega(\ell(Q_{k}(x)))}{\ell(Q_{k}(x))}\leq 2\int_{0}^{\ell(Q_{0})}\frac{\omega(t)}{t}\,\mathrm{d}t\lesssim[\omega]_{*}\omega(\ell(Q_{0})),$
and, since $\omega(\ell(Q_{0}))=\omega(|x-y|),$ we have shown that $\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n}[\omega]_{*}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
It is clear that the above argument gives $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$ when $\Gamma=\operatorname{small},$ as well as the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$ for $\Gamma=\operatorname{far}.$ The reverse inclusion ``""```$\supset$"` ` ⊃ " for $\Gamma=\operatorname{far}$ can be seen as follows. Let $R>0$ be such that if $\operatorname{\mathit{d}}(Q,0)>R,$ then $\mathcal{O}^{\omega}(f,Q)\leq\varepsilon.$ Let $|x|,|y|>R$ both be far from the origin. Using the structure of the Euclidean space, clearly there exists a dimensional constant $N(n)$ and points and cubes $\{z_{k},Q_{k}\}_{k=1}^{N}$ such that $z_{0}=x,$ $z_{N}=y$ and $z_{k},z_{k+1}\in Q_{k}$ and $\operatorname{\mathit{d}}(Q_{k},0)\geq R$ and $\ell(Q_{k})=|z_{k}-z_{k+1}|.$ Thus, by the chain of inequalities that led us to the right estimate for (29), we obtain
$|f(x)-f(y)|_{Y}\leq\sum_{k=1}^{N}|f(z_{k})-f(z_{k+1})|_{Y}\lesssim[\omega]_{*}\varepsilon\sum_{k=1}^{N}\omega(\ell(Q_{k}))\lesssim\varepsilon\omega(|x-y|),$
which concludes the proof of the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)\supset\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y).$
@@
$\lesssim_{C_{\operatorname{db}}}\Big{[}\varepsilon\Big{(}\frac{1}{\omega(2^{M+1})}\int_{2^{K}}^{2^{M+1}}\frac{\omega(t)}{t}\,\mathrm{d}t\Big{)}$
$\qquad\qquad\qquad+\frac{\omega(2^{K})}{\omega(2^{M+1})}\sup_{\ell(Q)\leq 2^{K}}\mathcal{O}^{\omega}(f;Q)\Big{(}\frac{1}{\omega(2^{K})}\int_{0}^{2^{K}}\frac{\omega(t)}{t}\,\mathrm{d}t\Big{)}\Big{]}\omega(\ell(Q_{0}))$
$\lesssim\Big{[}\varepsilon[\omega]_{*}+\frac{\omega(2^{K})}{\omega(2^{N+1})}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}[\omega]_{*}\Big{]}\omega(|x-y|).$
By the condition $\omega(\infty)=\infty$, we choose $N$ large enough and bound the bracketed term from above by $\leq 2\varepsilon[\omega]_{*}.$ Repeating the same proof with $y$ in place of $x$, we control the right sum on the line (29) and thus obtain $|f(x)-f(y)|\leq 4\varepsilon[\omega]_{*}\omega(|x-y|),$ that is, $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$
@@
$\leq\Big{[}\varepsilon\fint_{Q}\fint_{Q}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y+\sup_{\delta\leq R}\operatorname{osc}^{\omega}_{\delta}(f)\fint_{Q}\frac{1}{|Q|}\int_{Q_{R}(y)}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y\Big{]}|Q|.$
By repeating the bound (28), the left term in the square bracket is bounded from above by $\lesssim_{n,C_{\operatorname{db}}}\varepsilon\omega(\ell(Q)),$ which is a bound of the correct form. For the right term we bound
@@ 3. Approximation in infinite dimensional spaces
In this section we study smooth approximations in Banach spaces with respect to the $\dot{C}^{0,\omega}$ seminorm. In Subsections 3.1 and 3.2 we establish results for real-valued functions $f:X\to\mathbb{R}$, where the approximating functions are not only $C^{k}$ or $C^{1,\alpha}$ smooth but also Lipschitz and with bounded support. In the last Subsection 3.3 we obtain $C^{\infty}$ smooth approximations of Banach-valued mappings $f:c_{0}(\mathcal{A})\to Y,$ that are not necessarily Lipschitz, but have bounded support and belong to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ As a corollary, we deduce $C^{\infty}$ and $\dot{\operatorname{VC}}^{0,\omega}$ and approximations with bounded support for all $\dot{\operatorname{VC}}^{0,\omega}$ functions defined on $\mathbb{R}^{n}$ and with values in any Banach space.
@@
We will denote by $X^{*}$ the (continuous) dual of $X,$ and the dual norm in $X^{*}$ by $\|\cdot\|_{*},$ that is,
@@
[…]

### v4

@@ 2. Bounded support approximation
Let $f\in\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ and by Theorem 1.2 find $g\in\dot{\operatorname{VC}}_{\operatorname{small}}^{0,\omega}(\mathbb{R}^{n},Y)$ with bounded support such that $\|f-g\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}<\varepsilon.$ Let $r>0$ and $0\leq\eta_{r}\in C^{\infty}_{c}(B(0,r))$ be a standard real-valued smooth bump in $\mathbb{R}^{n}$ with $\int_{\mathbb{R}^{n}}\eta_{r}=1,$ and we define
$h_{r}(x)=g*\eta_{r}(x)=\int_{\mathbb{R}^{n}}g(x-z)\eta_{r}(z)\,\mathrm{d}z=\int_{B(0,r)}g(x-z)\eta_{r}(z)\,\mathrm{d}z,\quad x\in\mathbb{R}^{n},$
with the integral in the Bochner sense.
For the sake of completeness, let us justify why the integral $h_{r}(x)$ is well-defined for every $x\in\mathbb{R}^{n}.$ Denote momentarily $B:=B(0,r).$ The function
$B\ni z\mapsto\sigma_{x}(z):=g(x-z)\eta_{r}(z)$
is uniformly continuous $B\to Y$, by the compactness of $B$ and the continuity of the functions $g$ and $\eta_{r}.$ In particular, for each $y^{*}\in Y^{*},$ one has that the mapping $B\to\mathbb{R}$ given by $B\ni z\mapsto y^{*}(\sigma_{x}(z))$ is continuous, and thus Lebesgue measurable over $B.$ Also, the continuity of $\sigma_{x}$ implies that the image $\sigma_{x}(B)$ is compact, and so the closed linear span
$\overline{\mathrm{span}}\{\sigma_{x}(z)\,:\,z\in B\}$
is a separable subspace of $Y.$ By [BenyaminiLindenstrauss, Proposition 5.1], these two properties guarantee the measurability of $\sigma_{x}$ in $B.$ Then, by [BenyaminiLindenstrauss, Proposition 5.2] the Bochner-integrability of $\sigma_{x}$ is equivalent to $\|\sigma_{x}\|_{Y}\in L^{1}(B),$ that is, $\int_{B}\|\sigma_{x}(z)\|_{Y}\,\mathrm{d}z<\infty.$ But this is an immediate consequence of the fact that $\eta_{r}$ is a test function, and that $g,$ being a function of $\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)$, is bounded on bounded sets.
Notice that $h_{r}$ is compactly supported, as are $g$ and $\eta_{r}.$ Moreover, since $\eta_{r}\in C^{\infty}(\mathbb{R}^{n},Y)$ is with compact support, writing
$h_{r}(x)=\int_{\mathbb{R}^{n}}g(z)\eta_{r}(x-z)\,\mathrm{d}z,$
and using Differentiation Under the Integral Sign theorems for the Bochner integral (see [HajJoh14, Corollary 91]), one obtains that $h_{r}\in C^{\infty}(\mathbb{R}^{n},Y)$ as well.
Provided that $|u|\leq\delta,$ there holds uniformly in $r$ that
$|h_{r}(x+u)-h_{r}(x)|=\Big|\int_{\mathbb{R}^{n}}\eta_{r}(z)(g(x+u-z)-g(x-z))\,\mathrm{d}z\Big|\leq\|g\|_{\dot{C}^{0,\omega}_{\delta}}\omega(|u|),$
and hence we find $\delta>0$ so that, uniformly in $r,$ we have $\|g-h_{r}\|_{\dot{C}^{0,\omega}_{\delta}}\leq\|g\|_{\dot{C}^{0,\omega}_{\delta}}+\|h_{r}\|_{\dot{C}^{0,\omega}_{\delta}}\leq\varepsilon.$ To deal with the scales $\geq\delta$ we argue as follows. Observe that $g\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{small}}$ is uniformly continuous, by $\lim_{t\to 0}\omega(t)=0,$ and thus the approximation $h_{r}:=g*\eta_{r}$ converges uniformly to $g,$ as $r\to 0.$ We let $r=r(\delta)$ be so small that $\|g-h_{r}\|_{\infty}\leq\frac{1}{2}\varepsilon\omega(\delta).$ Then, for $|x-z|>\delta,$ there holds that
$\|(g-h_{r})(x)-(g-h_{r})(z)\|\leq 2\|g-h_{r}\|_{\infty}\leq\varepsilon\omega(\delta)\leq\varepsilon\omega(|x-z|),$
@@
For the reverse inclusion $\supset$", simply notice that by the Mean Value Inequality, $C^{\infty}_{c}(\mathbb{R}^{n},Y)\subset\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)$. We remind that this inequality holds for smooth $Y$-valued functions, as a consequence of the Hahn-Banach theorem on $Y.$ Also, by (3) there holds that $\operatorname{Lip}_{\operatorname{bs}}(\mathbb{R}^{n},Y)\subset\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y).$ As $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ is closed with respect to $\dot{C}^{0,\omega}$ limits, by Remark 2.1, we are done. ∎
@@
For a subset $G$ of $\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y),$ we write $G\subset_{u}\dot{\operatorname{VC}}^{0,\omega}(\mathbb{R}^{n},Y)$ (with u``u" for uniformly) provided that $G$ is equibounded in the $\dot{C}^{0,\omega}$ seminorm, and that for all $\varepsilon>0,$ there exists $t>0$ such that if |xy|<t|x-y|<t or $|x|>t^{-1}$ or $|y|>t^{-1},$ then
@@
We first prove the norms equivalent. Fix an arbitrary cube $Q$ and estimate
@@
For the other direction, fix $x,z\in\mathbb{R}^{n}$ and let $Q_{0}$ be a cube containing both $x$ and $y$ such that $\ell(Q_{0})=|x-y|.$ Consider the dyadic descendants of $Q_{0}$ (achieved by iteratively halving the sides) and for every integer $k\geq 0$ let $Q_{k}(x)$ be the descendant of $Q_{0}$ of sidelength $2^{-k}\ell(Q_{0})$ that contains the point $x.$ Similarly, we define $Q_{k}(y)$ for each $k.$ By the continuity of $f:\mathbb{R}^{n}\to Y,$ we may write
@@
Both sums are estimated identically, so it is enough to consider the first. We bound
@@
and we continue, using that $\omega$ is non-decreasing and the condition (7), with
$\sum_{k=0}^{\infty}\omega(\ell(Q_{k}(x)))=\sum_{k=0}^{\infty}\ell(Q_{k}(x))\frac{\omega(\ell(Q_{k}(x)))}{\ell(Q_{k}(x))}\leq 2\int_{0}^{\ell(Q_{0})}\frac{\omega(t)}{t}\,\mathrm{d}t\lesssim[\omega]_{*}\omega(\ell(Q_{0})).$
Since $\omega(\ell(Q_{0}))=\omega(|x-y|),$ it follows that $\|f\|_{\dot{C}^{0,\omega}(\mathbb{R}^{n},Y)}\lesssim_{n}[\omega]_{*}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}.$
It is clear that the above argument gives $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)=\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$, when $\Gamma=\operatorname{small},$ as well as the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\Gamma}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\Gamma}(\mathbb{R}^{n},Y)$, when $\Gamma=\operatorname{far}.$ The reverse inclusion $\supset$" for $\Gamma=\operatorname{far}$ can be seen as follows. Notice first that the assumption $f\in\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y),$ by the above proof, gives the following. For all $\varepsilon>0$ there exists some $M=M(\varepsilon)>0$ so that
$\sup_{\operatorname{\mathit{d}}(Q,0)\geq M}\sup_{x,y\in Q}\frac{|f(x)-f(y)|}{\omega(|x-y|)}\leq\varepsilon.$ (29)
If the line $[x,y]$ connecting $x$ to $y$ satisfies $\operatorname{\mathit{d}}([x,y],0)\geq M,$ then both $x,y$ can be enclosed in some large cube $Q$ such that $\operatorname{\mathit{d}}(Q,0)\geq M.$ Therefore, by (29), this case is clear.
Let $M=M(\varepsilon)$ be such that (29) holds and consider $x,y$ such that $|x|,|y|>R$ but $\operatorname{\mathit{d}}([x,y],0)<M$ and $R\gg M.$ We consider $M$ fixed and will show that taking $R=R(M,\|f\|_{\dot{C}^{0,\omega}},\omega)$ sufficiently large yields a correct bound. Since $R\gg M$, we find points $z_{1},z_{2}\in[x,y]$ such that $\operatorname{\mathit{d}}([x,z_{1}],0)\geq M$, $\operatorname{\mathit{d}}([z_{2},y],0)\geq M$ and $|z_{1}|,|z_{2}|\leq M.$ Then, using (29) for the pairs of points $(x,z_{1})$ and $(z_{2},y)$ we bound
$|f(x)-f(y)|\leq|f(x)-f(z_{1})|+|f(z_{1})-f(z_{2})|+|f(z_{2})-f(y)|\\ \leq\varepsilon\omega(|x-z_{1}|)+\|f\|_{\dot{C}^{0,\omega}}\omega(|z_{1}-z_{2}|)+\varepsilon\omega(|z_{2}-y|)\leq\Big(2\varepsilon+\frac{\omega(|z_{1}-z_{2}|)}{\omega(|x-y|)}\|f\|_{\dot{C}^{0,\omega}}\Big)\omega(|x-y|).$
To get a bound of a multiple of an $\varepsilon$ for the bracketed factor, we use that $|z_{1}-z_{2}|\leq|z_{1}|+|z_{2}|\leq 2M$ and that $|x-y|\geq R/2$. Indeed, for the latter bound, note that if |xy|<R/2|x-y|<R/2, then by $|x|,|y|>R$, for each $z\in[x,y]$, we can bound $|z|\geq|x|-|x-y|\geq R/2>M$ which contradicts $\operatorname{\mathit{d}}([x,y],0)\leq M.$ Using that $\omega$ is non-decreasing and $\omega(\infty)=\infty$ we take $R=R(M,\|f\|_{\dot{C}^{0,\omega}},\omega)$ so that
$\frac{\omega(|z_{1}-z_{2}|)}{\omega(|x-y|)}\|f\|_{\dot{C}^{0,\omega}}\leq\frac{\omega(4M)}{\omega(R)}\|f\|_{\dot{C}^{0,\omega}}\leq\varepsilon.$
This concludes the proof of the inclusion $\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y)\subset\mathrm{VMO}^{\omega}_{\operatorname{far}}(\mathbb{R}^{n},Y).$
@@
$\lesssim_{C_{\operatorname{db}}}\Big(\varepsilon\Big(\frac{1}{\omega(2^{M+1})}\int_{2^{K}}^{2^{M+1}}\frac{\omega(t)}{t}\,\mathrm{d}t\Big)$
$\qquad\qquad\qquad+\frac{\omega(2^{K})}{\omega(2^{M+1})}\sup_{\ell(Q)\leq 2^{K}}\mathcal{O}^{\omega}(f;Q)\Big(\frac{1}{\omega(2^{K})}\int_{0}^{2^{K}}\frac{\omega(t)}{t}\,\mathrm{d}t\Big)\Big)\omega(\ell(Q_{0}))$
$\lesssim\Big(\varepsilon[\omega]_{*}+\frac{\omega(2^{K})}{\omega(2^{N+1})}\|f\|_{\mathrm{BMO}^{\omega}(\mathbb{R}^{n},Y)}[\omega]_{*}\Big)\omega(|x-y|).$
By the condition $\omega(\infty)=\infty$, we choose $N$ large enough and bound the bracketed factor from above by $\leq 2\varepsilon[\omega]_{*}.$ Repeating the same proof with $y$ in place of $x$, we control the right sum on the line (28) and thus obtain $|f(x)-f(y)|\leq 4\varepsilon[\omega]_{*}\omega(|x-y|),$ that is, $f\in\dot{\operatorname{VC}}^{0,\omega}_{\operatorname{large}}(\mathbb{R}^{n},Y).$
@@
$\leq\Big(\varepsilon\fint_{Q}\fint_{Q}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y+\sup_{\delta\leq R}\operatorname{osc}^{\omega}_{\delta}(f)\fint_{Q}\frac{1}{|Q|}\int_{Q_{R}(y)}\omega(|x-y|)\,\mathrm{d}x\,\mathrm{d}y\Big)|Q|.$
By repeating the bound (27), the left term in the brackets is bounded from above by $\lesssim_{n,C_{\operatorname{db}}}\varepsilon\omega(\ell(Q)),$ which is a bound of the correct form. For the right term we bound
@@ 3. Approximation in infinite dimensional spaces
In this section we study approximations in Banach spaces with respect to the $\dot{C}^{0,\omega}$ seminorm. In Subsections 3.1 and 3.2 we establish results for real-valued functions $f:X\to\mathbb{R}$, where the approximating functions are not only $C^{k}$ or $C^{1,\alpha}$ smooth but also Lipschitz and with bounded support. In the last Subsection 3.3 we obtain $C^{\infty}$ smooth approximations of Banach-valued mappings $f:c_{0}(\mathcal{A})\to Y,$ that are not necessarily Lipschitz, but have bounded support and belong to $\dot{\operatorname{VC}}^{0,\omega}(X,Y).$ As a corollary, we deduce $C^{\infty}$ and $\dot{\operatorname{VC}}^{0,\omega}$ and approximations with bounded support for all $\dot{\operatorname{VC}}^{0,\omega}$ functions defined on $\mathbb{R}^{n}$ and with values in any Banach space.
@@
  • We denote by $X^{*}$ the (continuous) dual of $X,$ and the dual norm in $X^{*}$ by $\|\cdot\|_{*},$ that is,
@@
  • We say that a norm $\|\cdot\|$ on $X$ is of class $C^{k}$ if $\|\cdot\|\in C^{k}(X\setminus\{0\}),$ for $k\in\mathbb{N}\cup\{\infty\}.$